As a chemist for an agricultural products company, you have just developed a new herbicide,"Herbigon," that you think has the potential to kill weeds effectively. A sparingly soluble salt, Herbigon is dissolved in 1 MM acetic acid for technical reasons having to do with its production. You have determined that the solubility product Ksp of Herbigon is 8.50×10−6.

Although the formula of this new chemical is a trade secret, it can be revealed that the formula for Herbigon is X-acetate (XCH3COO, where "X" represents the top-secret cation of the salt). It is this cation that kills weeds. Since it is critical to have Herbigon dissolved (it won't kill weeds as a suspension), you are working on adjusting the pH so Herbigon will be soluble at the concentration needed to kill weeds. What pH must the solution have to yield a solution in which the concentration of X+ is 6.00×10−3 M? The pKa of acetic acid is 4.76.
Answer #1

pH = 1.91

Explanation

Given : concentration of X+ = 6.00 x 10-3 M

Ksp XCH3COO = 8.50 x 10-6

Ksp XCH3COO = [X+][CH3COO-]

8.50 x 10-6 = (6.00 x 10-3 M) * [CH3COO-]

[CH3COO-] = (8.50 x 10-6) / (6.00 x 10-3 M)

[CH3COO-] = 1.42 x 10-3 M

According to Henderson - Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log([CH3COO-] / [CH3COOH])

pH = 4.76 + log(1.42 x 10-3 M / 1 M)

pH = 4.76 + log(1.42 x 10-3)

pH = 4.76 + (-2.85)

pH = 1.91